Rationalize the denominator. 6/1-sqrt(5)
Answer by nshah11(47)

6/(1 - sqrt(5)) = 6(1 + sqrt(5))/(1 - sqrt(5))(1 + sqrt(5))
= (6 + 6sqrt(5))/(-4)
= -3/2*(1 + sqrt(5))